OpenStudy (adi3):

Will medal please help. consider g(n) = -12-2n/3. Find g^-1(n)

OpenStudy (adi3):

@misty1212

OpenStudy (adi3):

@ganeshie8

OpenStudy (adi3):

@Jhannybean

OpenStudy (jhannybean):

Is your function : \(g(n) = \dfrac{-12-2n}{3}\)?

OpenStudy (adi3):

yes

OpenStudy (adi3):

we need to find the inverse.

OpenStudy (jhannybean):

So first graph it to see whether it's a one-to-one function, and it is. \[y = \frac{-12-2n}{3}\]Switch y and n \[n=\frac{-12-2y}{3}\]Now lets solve for y.

OpenStudy (jhannybean):

I let my function \(g(n) = y\)

OpenStudy (jhannybean):

What function do you get when you resolve for y?

OpenStudy (adi3):

3n + 12 = -2y

OpenStudy (jhannybean):

Good, and when you isolate y?

OpenStudy (adi3):

that's where i reached till, i don't know what to do after that

OpenStudy (jhannybean):

To isolate the variable y, what do you have to do to each side of the equation?

OpenStudy (adi3):

divide by 2 i guess

OpenStudy (jhannybean):

-2, and yes, you were partially right.

OpenStudy (adi3):

so 3n + 12/-2 = y

OpenStudy (jhannybean):

(3n + 12)/ -2 = y * It helps when you put parenthesis () around the numerator and denominator

OpenStudy (adi3):

can I ask one more question please????

OpenStudy (jhannybean):

We can simplify that function a bit by dividing both terms in the numerator by -2. ((3n)/-2) +(12/-2) = y

OpenStudy (jhannybean):

Sure, I can try helping.

OpenStudy (adi3):

ok, find the inverse of the equation f(x) = \[-\frac{ 4 }{ 7 }x - \frac{ 16 }{ 7 }\]

OpenStudy (jhannybean):

\[y=\frac{-4x-16}{7}\]Same process, we switch x and y. \[x=\frac{-4y-16}{7}\] and now we resolve for y.

OpenStudy (adi3):

7x +16 = -4x

OpenStudy (adi3):

right?

OpenStudy (jhannybean):

-4y*

OpenStudy (adi3):

ohh yeah

OpenStudy (jhannybean):

7x + 16 = -4y and now we've got to isolate y, just like before.

OpenStudy (adi3):

\[\frac{ 7x+16 }{ -4 }\]

OpenStudy (jhannybean):

That's right, and after resolving for y, we relabel it as \(f^{-1}(x)\)

OpenStudy (adi3):

ok, can i ask further questions if i have doubt in them please?

OpenStudy (jhannybean):

Sure, but im logging off for a while, although good luck on learning this!`

OpenStudy (adi3):

when will you come back